# Taming Uncertainty in a Complex World: The Rise of Uncertainty Quantification — A Tutorial for Beginners

### v1

@@
George Box, a British statistician, wrote the famous aphorism, “All models are wrong, but some are useful.” The aphorism acknowledges that models, regardless of qualitative, quantitative, dynamical, or statistical, always fall short of the complexities of reality. The ubiquitous imperfections of models come from various sources, including the lack of a perfect understanding of nature, limited spatiotemporal model resolutions due to computational power, inaccuracy in the initial and boundary conditions, etc. Therefore, uncertainty quantification (UQ), which quantitatively characterizes and estimates uncertainties, is essential to identify the usefulness of the model. It also determines the output range if some aspects of the system are not precisely known. Exploring the uncertainty propagation, both through the time evolution and from one quantity to another via complicated nonlinear dependence, is one of the central goals of the UQ.
@@
Intuitively, the spread of a distribution, which describes how close the possible values of the output are to each other, measures the uncertainty of a variable. If a PDF is Gaussian, variance is a natural indicator describing the spread. Therefore, the uncertainty associated with the three Gaussian distributions is expected to increase from Panel (a) to Panel (c) in Figure 1. Despite this heuristic argument, there is a need to rigorously define and characterize the uncertainty. On the other hand, non-Gaussian distributions are widely seen in practice due to the intrinsic nonlinearity of the underlying system. The non-Gaussian features should also be highlighted in the rigorous quantification of the uncertainty.
@@
In the one-dimensional situation, (3) implies that the uncertainty is uniquely determined by the covariance and is independent of the mean value, consistent with the intuition. Therefore, Shannon’s entropy confirms that the uncertainty increases from Panel (a) to Panel (c) in Figure 1, and the uncertainty has the same value in Panel (a) and (d). However, there is typically no single indicator as the variance (or, equivalently, the peak height) that can be used to fully determine the uncertainty for complicated non-Gaussian PDFs. Therefore, Shannon’s entropy provides a systematic way for UQ. Panel (e) shows a Gamma distribution, which is skewed and has a one-sided fat tail. The fat tail usually corresponds to extreme events, which are farther from the mean value, naturally increasing the uncertainty. Therefore, despite the same peak height of the two PDFs in Panels (e) and (f), the fat-tailed Gamma PDF has a larger entropy.
@@
Figure 2: Solutions of the linear system (8). Panel (a): time evolution of $x(t)$ with a deterministic initial condition. Panel (b): time evolution of $x(t)$ with the initial condition given by a Gaussian distribution.
@@
Figure 3: Solutions of the nonlinear chaotic Lorenz 63 model (9). Panel (a): the Lorenz attractor. Panel (b): time evolution of $z(t)$ with a deterministic initial condition. Panel (b): time evolution of $z(t)$ with the initial condition given by a Gaussian distribution.
@@
which indicates that the higher-order moments containing the information of uncertainties affect the time evolution of the lower-order moments (e.g., the mean dynamics) via nonlinearity. It also reveals that the moment equations for general nonlinear dynamics will never be closed. In practice, approximations are made to handle the terms involving higher-order moments in the governing equations of the lower-order moments to form a solvable closed system.
@@
where $\propto$’ stands for ‘proportional to’. In (15), $p({u})$ is the forecast distribution by using a model built upon prior knowledge, while $p({v}|{u})$ is the probability of observation under the model assumption. Their combination is the conditional distribution of ${u}$ given $v$. Notably, due to the additional information from observation, uncertainty is expected to be reduced from the prior to the posterior distribution.
@@
where $I$ is an identity matrix of size $m\times m$ with $m$ being the dimension of $u$ and $K$ is given by $K=R_{f}g^{\mathtt{T}}(gR_{f}g^{\mathtt{T}}+{r}^{o})^{-1}$. Now consider the case with $m=1$ and $g=1$. Then, all three distributions in (15) are one-dimensional Gaussians, and the observation equals the truth plus noise. In such a case, $K=R_{f}/(R_{f}+r^{o})\in[0,1]$ and $\mu_{a}$ becomes a weighted summation of the prior mean and the observation with weights being $1-K$ and $K$, respectively. When the observational noise is much smaller than the prior uncertainty, i.e., $r^{o}\ll R_{f}$, the posterior mean almost fully trusts the observation since $K\approx 1$. In contrast, if the observation is highly polluted, i.e., $r^{o}\gg R_{f}$, then $K\approx 0$ and the observational information can almost be ignored, and the posterior mean nearly equals the prior mean. Essentially, the weights are fully determined by the uncertainties. On the other hand, the posterior variance $R_{a}$ in (16) is always no bigger than the prior variance $R_{f}$, indicating that the observation helps reduce the uncertainty. Panels (a)–(b) of Figure 4 validate the above conclusions. In some applications, repeated measurements with independent noises are available, which further advance the reduction of the posterior uncertainty. For example, in the above scalar state variable case, when $L$ repeated observations are used, the observational operator becomes a $L\times 1$ vector $g=(1,\ldots,1)^{\mathtt{T}}$ and similarly for the noise $\epsilon=(\epsilon_{1},\ldots,\epsilon_{L})^{\mathtt{T}}$. According to Panel (a), (c), and (d), the posterior uncertainty reduces when the number of observations increases from $L=1$ to $10$. Panel (e) shows that when $L$ becomes large, the posterior mean converges to the truth, and the posterior variance decreases to zero. Correspondingly, when characterizing the uncertainty reduction in the posterior distribution related to the prior via the relative entropy (7), the signal part converges to a constant while the dispersion part scales as $\ln(L)$.
Figure 4: Posterior distribution from Bayes formula (15) with $\mu_{f}=1$, $R_{f}=1$ and $m=1$. Panels (a)–(d): the PDFs with different observational uncertainty $r^{o}$ and number of observations $L$. Panel (e): the asymptotic behavior of the posterior mean $\mu_{a}$ and the posterior variance $R_{a}$ as a function of $L$. Due to the randomness in observations, the shading area shows the variation of the results with 100 sets of independent observations for each fixed $L$.
@@
Under certain conditions [chen2014information], closed analytic formulae are available for finding the posterior distribution of the nonlinear LaDA problem, which allows for using rigorous analysis to quantify the uncertainty reduction for the long-term behavior. By applying (7), it can be shown that the signal part of the relative entropy converges to a constant as $L\to\infty$, where the posterior mean $\mu$ in (7) is replaced by the true flow field. This is intuitive as the tracers become dense in the domain, and the observational error cancels off at each location when $L\to\infty$. On the other hand, the dispersion part will never converge. Instead, it grows at $\ln(L)$. Although it confirms that deploying extra tracers will always bring additional information, the uncertainty reduction will decrease as $L$ becomes large. Specifically, the logarithm dependence on $L$ implies that reducing the uncertainty by a fixed amount requires an exponential increase in the number of tracers. This means, in practice, once the information gain reaches a certain level, there is no need to continue deploying tracers as the uncertainty reduction from the additional tracers becomes marginal. Figure 5 shows a numerical simulation that validates the theoretic conclusion.
@@
The use of these eddy identification diagnostics requires knowing the exact flow field. However, uncertainties may appear in state estimation, for example, in the marginal ice zone when the ocean is estimated via the LaDA using a limited number of floe trajectories. Eddy identification exploiting the mean estimate of the flow field can lead to large biases.
@@
$\begin{split}\mathbb{E}[\mathrm{OW}(\mathbf{u})]=\mathrm{OW}(\bar{\mathbf{u}})&+\mathbb{E}[(u_{x}^{\prime})^{2}]-2\mathbb{E}[u_{x}^{\prime}v_{y}^{\prime}]\\ &+\mathbb{E}[(v_{y}^{\prime})^{2}]+4\mathbb{E}[v_{x}^{\prime}u_{y}^{\prime}].\end{split}$ (23)
@@
The governing equations of many complex turbulent systems are given by nonlinear partial differential equations (PDEs). The spectral method remains one of the primary schemes for finding numerical solutions. For simplicity, consider a scalar field $u(x,t)$, which satisfies a PDE with quadratic nonlinearities, as in many fluids and geophysical problems. Assume Fourier basis functions $\exp(ikx)$ are utilized, where $k$ is the wavenumber, and $x$ is the spatial coordinate. Denote by $\hat{u}_{k}(t)$ the time series of the spectral mode associated with wavenumber $k$. The resulting equation of $\hat{u}_{k}(t)$ typically has the following form,
@@
UQ plays a crucial role in calibrating the stochastic surrogate model. Specifically, the parameters in the stochastic surrogate model are optimized so that the two models have the same forecast uncertainty, which is crucial for DA and ensemble prediction. This is often achieved by matching a set of key statistics of the two models, especially the equilibrium PDFs and the decorrelation time. For (25), simple closed analytic formulae are available for model calibration in reproducing the forecast uncertainty, allowing it to be widely used in many practical problems.
@@
Panels (a) of Figure 7 shows a snapshot of the stream function from a two-layer quasi-geostrophic (QG) model, which is given by a set of nonlinear PDEs with the spatial resolution of each layer being $128\times 128$. In comparison, Panel (c) shows a snapshot of the spatial field by running a set of stochastic surrogate models, each calibrated by capturing the forecast uncertainty of one Fourier mode of the two-layer QG model [covington2022bridging]. Although the spatial resolution is only $23\times 23$, which contains less than 2% of the total number of modes, the spatial field is qualitatively similar to the QG model. Panels (b) and (d) show the time series of Fourier mode $(5,5)$ associated with the two models. They behave similarly as they have the same temporal correlation and PDFs.
@@ Conclusions
This paper exploits simple examples to provide the basic concepts of UQ. It highlights the fundamental changes in diagnostics when uncertainty appears. As uncertainty is ubiquitous, incorporating UQ into analysis or strategic planning is essential to facilitate understanding nature and problem-solving in almost all disciplines. UQ is not a stand-alone subject. It strongly relates to and advances the development of new techniques for mathematical reasoning, computational algorithms, effective modeling, data analysis, machine learning, and various applications. The areas in science, engineering, and technology requiring careful UQ are rapidly expanding and demands new ideas for coping with challenging issues.
@@
Supplementary Document and Codes Availability. This paper includes a supplementary document, which contains a comprehensive tutorial of many numerical examples shown in this article and beyond. The codes for these examples, written in both MATLAB and Python, are available from GitHub at https://github.com/marandmath/UQ_tutorial_code.

### v2

@@
George Box, a British statistician, wrote the famous aphorism, “All models are wrong, but some are useful.” The aphorism acknowledges that models, regardless of qualitative, quantitative, dynamical, or statistical, always fall short of the complexities of reality. The ubiquitous imperfections of models come from various sources, including the lack of a perfect understanding of nature, limited spatiotemporal model resolutions due to computational power, inaccuracy in the initial and boundary conditions, etc. Therefore, uncertainty quantification (UQ), which quantitatively characterizes and estimates uncertainties, is essential to identify the usefulness of the model. UQ also defines the range of possible outcomes when certain aspects of the system are not precisely known. A key objective of UQ is to explore how uncertainties propagate, both through time evolution and across different quantities via complex nonlinear dependencies.
@@
Intuitively, the spread of a distribution, which describes how close the possible values of the output are to each other, measures the uncertainty of a variable. If a PDF is Gaussian, variance is a natural indicator describing the spread. Therefore, the uncertainty associated with the three Gaussian distributions is expected to increase from Panel (a) to Panel (c) in Figure 1. On the other hand, non-Gaussian distributions are widely seen in practice due to the intrinsic nonlinearity of the underlying system. The non-Gaussian features should also be highlighted in the rigorous quantification of the uncertainty.
@@
In the one-dimensional situation, (3) implies that the uncertainty is uniquely determined by the covariance and is independent of the mean value, consistent with the intuition. Therefore, Shannon’s entropy confirms that the uncertainty increases from Panel (a) to Panel (c) in Figure 1, and the uncertainty has the same value in Panel (a) and (d). However, without a systematic information-based measurement, there is typically no single empirical indicator, like variance (or equivalently, peak height), that can fully characterize the uncertainty in complex non-Gaussian PDFs. Therefore, Shannon’s entropy provides a systematic way for UQ. Panel (e) shows a Gamma distribution, which is skewed and has a one-sided fat tail. The fat tail usually corresponds to extreme events, which are farther from the mean value, naturally increasing the uncertainty. Therefore, despite the same peak height of the two PDFs in Panels (e) and (f), the fat-tailed Gamma PDF has a larger entropy.
@@
Figure 2: Solutions of the linear system (8). Panel (a): time evolution of $x(t)$ with a deterministic initial condition. Panel (b): time evolution of $x(t)$ with the initial condition given by a Gaussian distribution. The ensemble size is 1000.
@@
Figure 3: Solutions of the nonlinear chaotic Lorenz 63 model (9). Panel (a): the Lorenz attractor. Panel (b): time evolution of $z(t)$ with a deterministic initial condition. Panel (b): time evolution of $z(t)$ with the initial condition given by a Gaussian distribution. The ensemble size is 1000.
@@
which indicates that the higher-order moments containing the information of uncertainties affect the time evolution of the lower-order moments (e.g., the mean dynamics) via nonlinearity. It also reveals that the moment equations for general nonlinear dynamics will never be closed. In practice, approximations are made to handle the terms involving higher-order moments in the governing equations of the lower-order moments to form a solvable closed system [majda2018strategies, chen2023stochastic].
@@
where $\propto$’ stands for ‘proportional to’. In (15), $p({u})$ is the forecast distribution by using a model built upon prior knowledge, while $p({v}|{u})$ is the probability of observation under the model assumption. Their combination is the conditional distribution of ${u}$ given $v$, which is called the posterior distribution. Notably, due to the additional information from observation, uncertainty is expected to be reduced from the prior to the posterior distribution.
@@
where $I$ is an identity matrix of size $m\times m$ with $m$ being the dimension of $u$ and $K$ is given by $K=R_{f}g^{\mathtt{T}}(gR_{f}g^{\mathtt{T}}+{r}^{o})^{-1}$. Now consider the case with $m=1$ and $g=1$. Then, all three distributions in (15) are one-dimensional Gaussians, and the observation equals the truth plus noise. In such a case, $K=R_{f}/(R_{f}+r^{o})\in[0,1]$ and $\mu_{a}$ becomes a weighted summation of the prior mean and the observation with weights being $1-K$ and $K$, respectively. When the observational noise is much smaller than the prior uncertainty, i.e., $r^{o}\ll R_{f}$, the posterior mean almost fully trusts the observation since $K\approx 1$. In contrast, if the observation is highly polluted, i.e., $r^{o}\gg R_{f}$, then $K\approx 0$ and the observational information can almost be ignored, and the posterior mean nearly equals the prior mean. Essentially, the weights are fully determined by the uncertainties. On the other hand, the posterior variance $R_{a}$ in (16) is always no bigger than the prior variance $R_{f}$, indicating that the observation helps reduce the uncertainty. Panels (a)–(b) of Figure 4 validate the above conclusions. In some applications, repeated measurements with independent noises are available, which further advance the reduction of the posterior uncertainty. For example, in the above scalar state variable case, when $L$ repeated observations are used, the observational operator becomes a $L\times 1$ vector $g=(1,\ldots,1)^{\mathtt{T}}$ and similarly for the noise $\epsilon=(\epsilon_{1},\ldots,\epsilon_{L})^{\mathtt{T}}$. According to Panel (a), (c), and (d), the posterior uncertainty reduces when the number of observations increases from $L=1$ to $10$. Panel (e) shows that when $L$ becomes large, the posterior mean converges to the truth, and the posterior variance decreases to zero. Correspondingly, when characterizing the uncertainty reduction in the posterior distribution related to the prior via the relative entropy (7), the signal part converges to a constant while the dispersion part scales as $\ln(L)$. See the online supplementary document for more details.
Figure 4: Posterior distribution from Bayes formula (15) with $\mu_{f}=1$, $R_{f}=1$ and $m=1$. Panels (a)–(d): the PDFs with different observational uncertainty $r^{o}$ and number of observations $L$. To keep the figure concise, the PDFs corresponding to the $L=5$ or $10$ noisy observations are omitted from Panels (c)–(d). Panel (e): the asymptotic behavior of the posterior mean $\mu_{a}$ and the posterior variance $R_{a}$ as a function of $L$. Due to the randomness in observations, the shading area shows the variation of the results with 100 sets of independent observations for each fixed $L$.
@@
Under certain conditions [chen2014information], closed analytic formulae are available for finding the posterior distribution of the nonlinear LaDA problem, which allows for using rigorous analysis to quantify the uncertainty reduction for the long-term behavior. By applying (7), it can be shown that the signal part of the relative entropy converges to a constant as $L\to\infty$, where the posterior mean $\mu$ in (7) is replaced by the true flow field. This is intuitive as the tracers become dense in the domain, and the observational error cancels off at each location when $L\to\infty$. On the other hand, the dispersion part in (7) will never converge. Instead, it grows at $\ln(L)$. Although it confirms that deploying extra tracers will always bring additional information, the increment of the uncertainty reduction will decrease as $L$ becomes large. Specifically, the logarithm dependence on $L$ implies that reducing the uncertainty by a fixed amount requires an exponential increase in the number of tracers. This means, in practice, once the information gain reaches a certain level, there is no need to continue deploying tracers as the uncertainty reduction from the additional tracers becomes marginal. Figure 5 shows a numerical simulation that validates the theoretic conclusion.
@@
The use of these eddy identification diagnostics requires knowing the exact flow field. However, uncertainties may appear in state estimation, for example, in the marginal ice zone when the ocean is estimated via the LaDA using a limited number of ice floe trajectories. Eddy identification exploiting the mean estimate of the flow field can lead to large biases.
@@
$\begin{split}\mathbb{E}[\mathrm{OW}(\mathbf{u})]=\mathrm{OW}(\bar{\mathbf{u}})&+\langle(u_{x}^{\prime})^{2}\rangle-2\langle u_{x}^{\prime}v_{y}^{\prime}\rangle\\ &+\langle(v_{y}^{\prime})^{2}\rangle+4\langle v_{x}^{\prime}u_{y}^{\prime}\rangle.\end{split}$ (23)
@@
The governing equations of many complex turbulent systems are given by nonlinear partial differential equations (PDEs). The spectral method remains one of the primary schemes for finding numerical solutions. For simplicity, consider a scalar field $u(x,t)$, which satisfies a PDE with quadratic nonlinearities, as in many fluids and geophysical problems. Assume Fourier basis functions $\exp(ikx)$ are utilized, where $k$ is the wavenumber, and $x$ is the spatial coordinate. Denote by $\hat{u}_{k}(t)$ the time series of the spectral mode associated with wavenumber $k$. Apply a finite-dimensional truncation to retain only the modes $k\in[-K,K]$. The resulting equation of $\hat{u}_{k}(t)$ typically has the following form,
@@
UQ plays a crucial role in calibrating the stochastic surrogate model. Specifically, the parameters in the stochastic surrogate model are optimized so that the two models have the same forecast uncertainty, which is crucial for DA and ensemble prediction. This is often achieved by matching a set of key statistics of the two models, especially the equilibrium PDFs and the decorrelation time. For (25), simple closed analytic formulae are available for model calibration in reproducing the forecast uncertainty, allowing it to be widely used in many practical problems. See the online supplementary document for more details.
@@
Panels (a) of Figure 7 shows a snapshot of the stream function from a two-layer quasi-geostrophic (QG) model, which is given by a set of nonlinear PDEs with the spatial resolution of each layer being $128\times 128$. In comparison, Panel (c) shows a snapshot of the spatial field by running a set of stochastic surrogate models, each calibrated by capturing the forecast uncertainty of one Fourier mode of the two-layer QG model [covington2022bridging]. Although the spatial resolution is only $23\times 23$, which contains less than 2% of the total number of modes, the spatial field is qualitatively similar to the QG model. Panels (b) and (d) show the time series of Fourier mode $(5,5)$ associated with the two models. They behave similarly as they have the same temporal correlation and PDFs. An additional example of using the stochastic model (25) to approximate the KdV equation can be found in [majda2019statistical].
@@ Conclusions
This paper exploits simple examples to provide the basic concepts of UQ. It does not aim to provide a single, comprehensive definition, as various communities currently have different interpretations of UQ. Nevertheless, the information quantities (1) and (6) are natural measurements to quantify the uncertainty. As a developing field within applied mathematics and interdisciplinary research, UQ continues to evolve. One of the goals of this paper is to offer new insights into how these ideas can be applied across different fields, helping to reveal the commonalities and practical advantages of diverse approaches. As uncertainty is ubiquitous, incorporating UQ into analysis or strategic planning is essential to facilitate understanding nature and problem-solving in almost all disciplines. UQ has become an important tool in tackling observational data [braverman2021post], forecasting turbulent systems [majda2012lessons], and in studying inverse problems [dashti2011uncertainty]. UQ is not a stand-alone subject. It strongly relates to and advances the development of new techniques for mathematical reasoning, computational algorithms, effective modeling, data analysis, machine learning, and various applications. The areas in science, engineering, and technology requiring careful UQ are rapidly expanding and demands new ideas for coping with challenging issues.
@@
Supplementary Document and Codes Availability. The arXiv version of this article includes a supplementary document (https://arxiv.org/abs/2408.01823), which contains a comprehensive tutorial of many numerical examples shown in this article and beyond. The codes for these examples, written in both MATLAB and Python, are available from GitHub at https://github.com/marandmath/UQ_tutorial_code.